# Variational MineGAN: A Data-efficient Knowledge Transfer Architecture for Generative AI-assisted Design of Nanophotonic Structures

### v1

@@
Abstract
Leveraging the power of deep learning to design nanophotonic devices has been an area of active research in recent times, with Generative Adversarial Networks (GANs) being a popular choice alongside autoencoder-based methods. However, both approaches typically require large datasets and significant computational resources, which can outweigh the advantages of saving time and effort. While GANs trained on smaller datasets can experience challenges such as unstable training and mode collapse, fine-tuning pre-trained GANs on these limited datasets often introduces additional problems, such as overfitting and lack of flexibility. This limits the model’s ability to generalize, reducing its overall effectiveness. In this study, we present Variational MineGAN, an enhanced version of MineGAN that is less susceptible to overfitting while leveraging pre-trained GANs. This approach ensures a more robust sampling process, improving the model’s generalization ability while transferring knowledge to domains with limited data. Experimental results demonstrate a lower Frechet Inception Distance (FID) score of 52.14, along with an increased Inception Score (IS) of 3.59 compared to prior methods. This implies better quality design images, which allow for the exploration of highly efficient designs with desired spectral responses and improved learning of nonlinear relationships between the latent space and the corresponding designs.
keywords:
Photonics, Metamaterial, Generative Adversarial Network, Transfer learning, Deep learning, Inverse design
@@ Abstract
{tocentry}
@@ 1 Introduction
Understanding the impact of intricate geometrical shapes on the properties of micro or nanophotonic devices has been among the key challenges in photonics research1, 2. The conventional design approach typically begins with an educated guess about the structure’s shape, followed by the use of Finite-Difference Time-Domain (FDTD) or Finite Element Method (FEM) solver software to solve Maxwell’s partial differential equations 3, 4to obtain the spectral response of the structure. The design parameters are then tweaked to optimize performance and iteratively evaluated through computationally expensive simulations. This trial-and-error method is both time and resource-intensive 5, 6, 7.
@@
To address these limitations, different methods based on deep learning (DL) have recently been employed for the design of nanophotonic structures. Discriminative deep learning models have been utilized for both forward and inverse electrodynamics problems 8, 9, 10. Additionally, generative deep learning models such as Generative Adversarial Networks (GANs)11 and Variational Autoencoders (VAEs)12 have been utilized for design optimization. Different variations of generative model architectures have been proven to be effective in solving these optimization problems. Kudyshev et al. used an adversarial autoencoder for topology optimization of thermal emitters 13. Panisilvam et al. proposed the use of asymmetric CycleGAN for the inverse design of photonic devices 14. These generative models excel at learning the underlying distribution of data and generating realistic images, making them a popular choice for optimizing nanophotonic structures based on desired spectral responses. However, they typically require large datasets and extensive training time, which can be a significant limitation.15, 16
@@
To mitigate these challenges, the concept of “knowledge transfer” has been applied in generative model training. This approach leverages pre-trained generative models trained on large datasets for a longer time and adapts them to new, smaller datasets. Weng et al. demonstrated the effectiveness of using knowledge transfer to fine-tune the pre-trained GAN model on limited data 17. However, Noguchi and Harada addressed the issue of mode collapse associated with this approach and proposed freezing certain parameters of the generator18. While this approach is less prone to overfitting, it significantly limits the model’s flexibility during fine-tuning19.
@@
To further address the shortcomings of transfer learning of the GAN model on limited data, Wang et al. introduced the “MineGAN” approach 19. This method employs a miner network that transforms the input to a new distribution of target data, which facilitates the generator model in generating images from the new target domain.
@@
In this work, we introduce “Variational MineGAN,” an extended version of MineGAN. Our approach modifies the miner network to directly learn the properties of the target data distribution by estimating its mean and variance instead of transforming the input to follow that distribution. This allows better understanding and control over the latent distribution space and reduces computational cost as well since it bypasses the use of the miner network during inference, streamlining the image generation process.
@@ 2.1 GAN Formulation
A Generative Adversarial Network (GAN) comprises a pair of networks: the Generator and the Discriminator. The discriminator’s job is to differentiate between actual and generated fake samples, while the generator’s job is to map random noise samples into images. Goodfellow et al. presented the initial GAN formulation 20, which seeks to solve a two-player min-max problem by reaching a Nash equilibrium, which is mathematically expressed as:
@@
in which $z\epsilon R^{d_{z}}$ is a latent variable obtained from a distribution $p(z)$ typically a normal distribution with identity covariance $\mathcal{N}(0,I)$ or uniform distribution $\mathcal{U}[-1,1]$. On the other hand, $x\sim p_{\textnormal{data}}(x)$ corresponds to samples drawn from the real data distribution. In this work, for better training stability, Wasserstein GAN-Gradient Penatly (WGAN-GP) 21 loss has been used for the discriminator, which utilizes the Wasserstein Loss 22 alongside a penalty term added to it, instead of Binary Cross Entropy (BCE) Loss. The discriminator’s modified loss function then takes on the following form:
@@
$\hat{x}=\alpha x_{real}+(1-\alpha)x_{fake}$ (3)
where $\alpha$ sampled from a uniform distribution $\mathcal{U}[-1,1]$ 21.
Figure 1: Overview of the ProGAN architecture. (a) ProGAN Generator: progressively generates 4x4 to 64x64 images from input noise. (b) Building blocks used in generator and discriminator. (c) ProGAN Discriminator: makes a real/fake prediction from input image samples.
In this work, for both the generator and discriminator, we follow the Progressive GAN (Pro-GAN) architecture 14, where the training begins with images of lower resolution, and the resolution is progressively enhanced by including additional layers to the network. Although the original Pro-GAN paper starts training at 4$\times$4 and goes up to 1024$\times$1024 resolution, due to dataset constraints, we train up to 64$\times$64 resolution images. The latent noise dimension used in this work is 256, and the noise vector is drawn from $\mathcal{N}(0,I)$. The generator and discriminator network architecture has been illustrated in Figure 1. The convolution layers inside the “Convolution Block” (Conv. Block) employ a kernel size of 3, stride of 1, and padding of 1. The generator uses 256 filters for convolution blocks at 4$\times$4 to 32$\times$32 image sizes and 128 filters for the 64$\times$64 size. In contrast, the discriminator follows the opposite filter setup. In the discriminator’s final block, the three convolution layers use kernel sizes of 3, 4, and 1 with no padding. The output of the last convolution layer skips activation, as WGAN is used. The value of $\lambda$ for WGAN-GP is set to 10. The core of the generator and the discriminator network are mirror images of each other, and they both always grow in harmony. This ProGAN generator and discriminator are later on used as pre-trained models for both MineGAN and Variational MineGAN.
@@ 2.2 MineGAN
MineGAN employs a miner network that steers the input noise to the most promising regions of the latent space with respect to the target data distribution $D^{T}$19. The miner network is a compact neural network composed of a sequence of linear layers, each preceding ReLU activation and batch normalization except for the last layer. Each linear layer has a dimension of 256. Figure 2 (a) illustrates the detailed configuration of the linear blocks used in the miner network and the configuration of the miner network itself used in our implementation.
@@
MineGAN training takes place in two stages. Initially, the generator weights are kept fixed, and only the miner is trained. In this stage, the miner network learns to transform the input noise $z\sim p(z)$ in such a manner that it follows the new target data distribution. The generator is no longer fixed during the second stage of training, effectively fine-tuning the entire architecture. However, since the miner network is composed of significantly less amount of parameters compared to the GAN network, it is, therefore, less susceptible to overfitting19. It narrows the latent distribution to align more with the target, simplifying fine-tuning by providing a more consistent training signal, and as a result, the generator weights do not require to go through significant updates, thus reducing the likelihood of overfitting. The employed loss functions for the Generator and the Discriminator are
@@ 2.4 Absorption Spectra Prediction Model
In addition to training the generative models, a separate predictive model was also trained to estimate the absorption spectra based on the design images generated. A Convolutional Neural Network (CNN) makes up this model that takes a 64$\times$64 pixel grayscale design image as input, and 80 spectral values corresponding to wavelengths ranging from 4 nm to 12 nm come out as output. The output of this model is passed through a smoothing filter with a window size of 5 to obtain a smooth spectral response. The architecture of this model has been illustrated in Figure 4. This model allows us to efficiently select designs that meet specific criteria before conducting more resource-intensive simulations for further validation.
@@ 2.5 Dataset Details
To pre-train our ProGAN model, we utilized a publicly available dataset from Raman Lab of UCLA comprising 18,770 images of metamaterial absorbers 23. Lumerical software’s 3D Finite-Difference Time-Domain (FDTD) simulations were used to create these data. The dataset features seven distinct shapes, each with geometric parameter variations. These structures had 3.2$\times$3.2 $\mu m^{2}$ as their unit cell dimensions. In order to simulate periodic structures, perfectly matched layers (PMLs) were employed above and below the plane of interest, and periodic boundary conditions were implemented on the side boundaries. A source plane was injected parallel to and at a distance from the plane of interest. Reflectance and transmission spectra were calculated using the above-mentioned simulation arrangement.
For the fine-tuning stage with MineGAN and Variational MineGAN, we constructed a smaller target dataset of 1000 images by combining images from another dataset from Raman lab 24 with an additional 200 images of designs that we simulated, which involved variations of different geometric structures. The physical configuration of the designs from this dataset is similar to the one described previously. This dataset also contains the absorption spectra within the range of 4nm to 12nm wavelength on a separate file. This dataset was used to train the MineGAN and Variational MineGAN models, as well as the spectra prediction model, incorporating a diverse range of geometric shapes and sizes. This approach ensured that the target dataset introduced new variations that the pre-trained model had not encountered before.
@@ 2.6.1 ProGAN pre-training
For the pre-trained network, a ProGAN architecture was trained on the multiclass dataset of Raman Lab 23. Training started at 4$\times$4 resolution, then 8x8, 16$\times$16 and gradually up to 64$\times$64 pixels. The models were trained for 30 epochs at each resolution. PixelNorm was applied after each convolution layer, and minibatch standard deviation information was integrated into the channel dimension of the discriminator. The learning rate was configured to 0.001 using the Adam optimizer, Lambda-GP was set to 10, and one discriminator iteration was performed per generator update.
@@ 3 Results and Discussion
We conducted a comprehensive evaluation of four different generative approaches: training from scratch (ProGAN trained directly on the target dataset), Transfer GAN 17, MineGAN, and our proposed Variational MineGAN. Each model was evaluated using two key performance metrics, Frechet Inception Distance (FID)25 and Inception Score (IS) 26. FID measures the similarity between real and generated images by comparing the distributions of feature representations. On the other hand, IS evaluates the quality and diversity of generated images by assessing how confidently the images are classified into distinct classes. A lower FID score and higher IS are desirable since they indicate that the generated images are realistic and exhibit a greater diversity of features27, 28.
@@
The FID scores, in particular, show a marked difference when compared to the models trained from scratch or using Transfer GAN without a miner. The variational component not only improves the model’s generalization capabilities but also reduces overfitting, which is a common challenge in GAN training, particularly with small datasets29. Figure 5 shows that Variational MineGAN generates clearer, more realistic images compared to other approaches. Images from models trained from scratch and Transfer GAN show more artifacts and lack sharpness, which becomes more prominent in smaller datasets.
@@
To further investigate the degree of overfitting, we conducted an interpolation experiment by generating images from interpolations between two latent noise vectors. This test helps to assess how well each model captures the global structure of the data. Models that overfit tend to generate disjointed or abrupt transitions between interpolated samples, while models that generalize well will produce smooth and coherent transitions30, 31. As illustrated in figure 7, Variational MineGAN produces smooth and continuous image transitions, demonstrating its ability to identify the underlying data distribution effectively. On the other hand, Transfer GAN produced more erratic transitions, indicating a higher level of overfitting to the training data. This observation further supports the conclusion that Variational MineGAN better generalizes to unseen data.
@@
Figure 8: Real and predicted spectra values for specific designs, with the corresponding design images and MSE values displayed as insets.
@@ 5 Data Availability
6 Acknowledgement
The authors acknowledge that no external funding was provided for this study.

### v2

@@
Variational MineGAN: A Data-efficient Knowledge Transfer Architecture for Generative AI-assisted Design of Nanophotonic Structures
@@
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Using deep learning to design nanophotonic devices has recently been an active research area, with Generative Adversarial Networks (GANs) being a popular choice alongside autoencoder-based methods. However, both generally require large datasets and computational power, creating limitations in data-scarce scenarios. Fine-tuning GANs on limited data often cause mode collapse and overfitting, reducing generalizability and model effectiveness. To address this, we introduce Variational MineGAN, an optimized architecture that enhances data efficiency by reducing overfitting. Experimental results demonstrate a Frechet Inception Distance (FID) of 52.14 and an Inception Score (IS) of 3.59, enabling high-quality design generation and accurate spectral response estimation, improving nanophotonics design exploration.
@@ 1 Introduction
Understanding the impact of intricate geometrical shapes on the properties of micro or nanophotonic devices has been among the key challenges in photonics research[1, 2]. The conventional design approach typically begins with an educated guess about the structure’s shape, followed by the use of Finite-Difference Time-Domain (FDTD) or Finite Element Method (FEM) solver software to solve Maxwell’s partial differential equations [3, 4]to obtain the spectral response of the structure. The design parameters are then tweaked to optimize performance and iteratively evaluated through computationally expensive simulations. This trial-and-error method is both time and resource-intensive [5, 6, 7].
@@
To address these limitations, different methods based on deep learning (DL) have recently been employed for the design of nanophotonic structures. Discriminative deep learning models have been utilized for both forward and inverse electrodynamics problems [8, 9, 10]. Additionally, generative deep learning models such as Generative Adversarial Networks (GANs)[11] and Variational Autoencoders (VAEs)[12] have been utilized for design optimization. Different variations of generative model architectures have been proven to be effective in solving these optimization problems. Kudyshev et al. used an adversarial autoencoder for topology optimization of thermal emitters [13]. Panisilvam et al. proposed the use of asymmetric CycleGAN for the inverse design of photonic devices [14]. These generative models excel at learning the underlying distribution of data and generating realistic images, making them a popular choice for optimizing nanophotonic structures based on desired spectral responses. However, they typically require large datasets and extensive training time, which can be a significant limitation.[15, 16]
@@
To mitigate these challenges, the concept of “knowledge transfer" has been applied in generative model training. This approach leverages pre-trained generative models trained on large datasets for a longer time and adapts them to new, smaller datasets. Weng et al. demonstrated the effectiveness of using knowledge transfer to fine-tune the pre-trained GAN model on limited data [17]. However, Noguchi and Harada addressed the issue of mode collapse associated with this approach and proposed freezing certain parameters of the generator[18]. While this approach is less prone to overfitting, it significantly limits the model’s flexibility during fine-tuning[19].
@@
To further address the shortcomings of transfer learning of the GAN model on limited data, Wang et al. introduced the “MineGAN" approach [19]. This method employs a miner network that transforms the input to a new distribution of target data, which facilitates the generator model in generating images from the new target domain.
@@
In this work, we introduce “Variational MineGAN," an extended version of MineGAN. Our approach modifies the miner network to directly learn the properties of the target data distribution by estimating its mean and variance instead of transforming the input to follow that distribution. This allows better understanding and control over the latent distribution space and reduces computational cost as well since it bypasses the use of the miner network during inference, streamlining the image generation process.
@@ 2.1 GAN Formulation
A Generative Adversarial Network (GAN) comprises a pair of networks: the Generator and the Discriminator. The discriminator’s job is to differentiate between actual and generated fake samples, while the generator’s job is to map random noise samples into images. Goodfellow et al. presented the initial GAN formulation [20], which seeks to solve a two-player min-max problem by reaching a Nash equilibrium, which is mathematically expressed as:
@@
in which $z\epsilon R^{d_{z}}$ is a latent variable obtained from a distribution $p(z)$ typically a normal distribution with identity covariance $\mathcal{N}(0,I)$ or uniform distribution $\mathcal{U}[-1,1]$. On the other hand, $x\sim p_{\textnormal{data}}(x)$ corresponds to samples drawn from the real data distribution. In this work, for better training stability, Wasserstein GAN-Gradient Penatly (WGAN-GP) [21] loss has been used for the discriminator, which utilizes the Wasserstein Loss [22] alongside a penalty term added to it, instead of Binary Cross Entropy (BCE) Loss. The discriminator’s modified loss function then takes on the following form:
@@
$\hat{x}=\alpha x_{real}+(1-\alpha)x_{fake},$ (3)
where $\alpha$ is sampled from a uniform distribution $\mathcal{U}[-1,1]$ [21].
Figure 1: Overview of the ProGAN architecture. (a) ProGAN Generator: progressively generates 4$\times$4 to 64$\times$64 images from input noise. (b) Building blocks used in generator and discriminator. (c) ProGAN Discriminator: makes a real/fake prediction from input image samples.
In this work, for both the generator and discriminator, we follow the Progressive GAN (Pro-GAN) architecture [14], where the training begins with images of lower resolution, and the resolution is progressively enhanced by including additional layers to the network. Although the original Pro-GAN paper starts training at 4$\times$4 and goes up to 1024$\times$1024 resolution, due to dataset constraints, we train up to 64$\times$64 resolution images. The latent noise dimension used in this work is 256, and the noise vector is drawn from $\mathcal{N}(0,I)$. The generator and discriminator network architecture has been illustrated in Figure 1. The convolution layers inside the “Convolution Block" (Conv. Block) employ a kernel size of 3, stride of 1, and padding of 1. The generator uses 256 filters for convolution blocks at 4$\times$4 to 32$\times$32 image sizes and 128 filters for the 64$\times$64 size. In contrast, the discriminator follows the opposite filter setup. In the discriminator’s final block, the three convolution layers use kernel sizes of 3, 4, and 1 with no padding. The output of the last convolution layer skips activation, as WGAN is used. The value of $\lambda$ for WGAN-GP is set to 10. The core of the generator and the discriminator network are mirror images of each other, and they both always grow in harmony. This ProGAN generator and discriminator are later on used as pre-trained models for both MineGAN and Variational MineGAN.
@@ 2.2 MineGAN
MineGAN employs a miner network that steers the input noise to the most promising regions of the latent space with respect to the target data distribution $D^{T}$[19]. The miner network is a compact neural network composed of a sequence of linear layers, each preceding ReLU activation and batch normalization except for the last layer. Each linear layer has a dimension of 256. Figure 2 (a) illustrates the detailed configuration of the linear blocks used in the miner network and the configuration of the miner network itself used in our implementation.
@@
MineGAN training takes place in two stages. Initially, the generator weights are kept fixed, and only the miner is trained. In this stage, the miner network learns to transform the input noise $z\sim p(z)$ in such a manner that it follows the new target data distribution. The generator is no longer fixed during the second stage of training, effectively fine-tuning the entire architecture. However, since the miner network is composed of significantly less amount of parameters compared to the GAN network, it is, therefore, less susceptible to overfitting[19]. It narrows the latent distribution to align more with the target, simplifying fine-tuning by providing a more consistent training signal, and as a result, the generator weights do not require to go through significant updates, thus reducing the likelihood of overfitting. The employed loss functions for the Generator and the Discriminator are
@@ 2.4 Absorption Spectra Prediction Model
[…]
@@ 2.5 Dataset Details
To pre-train our ProGAN model, we utilized a publicly available dataset from Raman Lab of UCLA comprising 18,770 images of metamaterial absorbers [23]. Lumerical software’s 3D Finite-Difference Time-Domain (FDTD) simulations were used to create these data. The dataset features seven distinct shapes, each with geometric parameter variations. These structures had 3.2$\times$3.2 $\mu m^{2}$ as their unit cell dimensions. In order to simulate periodic structures, perfectly matched layers (PMLs) were employed above and below the plane of interest, and periodic boundary conditions were implemented on the side boundaries. A source plane was injected parallel to and at a distance from the plane of interest. Reflectance and transmission spectra were calculated using the above-mentioned simulation arrangement.
[…]
@@ 2.6.1 ProGAN pre-training
For the pre-trained network, a ProGAN architecture was trained on the multiclass dataset of Raman Lab [23]. Training started at 4$\times$4 resolution, then 8x8, 16$\times$16 and gradually up to 64$\times$64 pixels. The models were trained for 30 epochs at each resolution. PixelNorm was applied after each convolution layer, and minibatch standard deviation information was integrated into the channel dimension of the discriminator. The learning rate was configured to 0.001 using the Adam optimizer, Lambda-GP was set to 10, and one discriminator iteration was performed per generator update.
@@ 3 Results and Discussion
We conducted a comprehensive evaluation of four different generative approaches: training from scratch (ProGAN trained directly on the target dataset), Transfer GAN [17], MineGAN, and our proposed Variational MineGAN. Each model was evaluated using two key performance metrics, Frechet Inception Distance (FID)[25] and Inception Score (IS) [26]. FID measures the similarity between real and generated images by comparing the distributions of feature representations. On the other hand, IS evaluates the quality and diversity of generated images by assessing how confidently the images are classified into distinct classes. A lower FID score and higher IS are desirable since they indicate that the generated images are realistic and exhibit a greater diversity of features[27, 28].
@@
The FID scores, in particular, show a marked difference when compared to the models trained from scratch or using Transfer GAN without a miner. The variational component not only improves the model’s generalization capabilities but also reduces overfitting, which is a common challenge in GAN training, particularly with small datasets[29]. Figure 5 shows that Variational MineGAN generates clearer, more realistic images compared to other approaches. Images from models trained from scratch and Transfer GAN show more artifacts and lack sharpness, which becomes more prominent in smaller datasets.
@@
To further investigate the degree of overfitting, we conducted an interpolation experiment by generating images from interpolations between two latent noise vectors. This test helps to assess how well each model captures the global structure of the data. Models that overfit tend to generate disjointed or abrupt transitions between interpolated samples, while models that generalize well will produce smooth and coherent transitions[30, 31]. As illustrated in figure 7, Variational MineGAN produces smooth and continuous image transitions, demonstrating its ability to identify the underlying data distribution effectively. On the other hand, Transfer GAN produced more erratic transitions, indicating a higher level of overfitting to the training data. This observation further supports the conclusion that Variational MineGAN better generalizes to unseen data.
@@
Figure 8: Real and predicted spectra values for specific generated designs, with the corresponding design images, simulated versions and MSE values displayed as insets.
@@ 4 Conclusion
Funding No external funding was available for this research.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability The code, along with the FDTD simulation data used for this work, are available at the following link: https://github.com/Tarvirator/Variational-MineGAN